# Reason to Rote: Rethinking Memorization in Reasoning

### v1

@@ Abstract
Yupei Du1, Philipp Mondorf2, Silvia Casola2, Yuekun Yao3, Robert Litschko2, and Barbara Plank2 1 Department of ICS, Utrecht University, the Netherlands 2 MaiNLP, Center for Information and Language Processing, LMU Munich, Germany and Munich Center for Machine Learning (MCML) 3 Saarland Informatics Campus, Saarland University, Saarbrücken, Germany 1y.du@uu.nl, 2{p.mondorf,s.casola,robert.litschko,b.plank}@lmu.de, 3ykyao@coli.uni-saarland.de
@@ 2.1 Tasks
Four-Digit Addition
@@ 5 Related work
LLMs are well-known to memorize verbatim, which may pose privacy and copyright concerns (Lukas et al., 2023; Karamolegkou et al., 2023). For example, Carlini et al. (2021) show that hundreds of training examples can be extracted from GPT-2; Carlini et al. (2023) studies the factors that influence memorization, including model capacity, number of duplications, and prompt context length. Moreover, Biderman et al. (2023) show such memorization is predictable. The most closely related work is Huang et al. (2024), which studies verbatim memorization of training data in LLMs. They show that memorization is distributed across tokens and builds on the model’s general LM capabilities. However, their work does not explore the co-existence of generalization and memorization; moreover, it does not examine the specific mechanism of memorization from intermediate steps.
@@ 6 Discussion
We performed most of our experiments on synthetic datasets. While being artificially generated, they represent real-world tasks: namely arithmetic addition and question answering based on multi-hop relational reasoning. In other words, rather than learning imaginary tasks, models learn to solve real-world problems with carefully controlled settings, enabling fine-grained analyses of model behavior. Previous studies have demonstrated that this approach yields deep insights into language models and produces impactful, practical findings (Power et al., 2022; Pearce et al., 2023; Allen-Zhu and Li, 2024; Wang et al., 2024a).

### v2

@@ Abstract
Yupei Du1††thanks: Work done partly during a research visit of YD to LMU Munich, which is supported by Utrecht University. YD is now affiliated with Saarland University. , Philipp Mondorf2, Silvia Casola2, Yuekun Yao3, Robert Litschko2, and Barbara Plank2 1 Department of ICS, Utrecht University, the Netherlands 2 MaiNLP, Center for Information and Language Processing, LMU Munich, Germany and Munich Center for Machine Learning (MCML) 3 Saarland Informatics Campus, Saarland University, Saarbrücken, Germany 1y.du@uu.nl, 2{p.mondorf,s.casola,robert.litschko,b.plank}@lmu.de, 3ykyao@coli.uni-saarland.de
@@ 2.1 Tasks
Four‑Digit Addition
@@ 5 Related work
LLMs are well-known to memorize verbatim, which may pose privacy and copyright concerns (Lukas et al., 2023; Karamolegkou et al., 2023). For example, Carlini et al. (2021) show that hundreds of training examples can be extracted from GPT-2; Carlini et al. (2023) studies the factors that influence memorization, including model capacity, number of duplications, and prompt context length. Moreover, Biderman et al. (2023) show such memorization is predictable. However, recent work by Liu et al. (2025) show that the commonly-used completion test, i.e., n-gram based membership inference, is not reliable, by producing verbatim texts that are not part of the training data. The most closely related work is Huang et al. (2024), which studies verbatim memorization of training data in LLMs. They show that memorization is distributed across tokens and builds on the model’s general LM capabilities. However, their work does not explore the co-existence of generalization and memorization; moreover, it does not examine the specific mechanism of memorization from intermediate steps.
@@ 6 Discussion
We performed most of our experiments on synthetic datasets, which, though artificially generated, capture real-world tasks such as arithmetic addition and multi-hop relational reasoning. In other words, rather than learning imaginary tasks, models learn to solve real-world problems with carefully controlled settings, enabling fine-grained analyses of model behavior. Previous studies have demonstrated that this approach yields deep insights into language models and produces impactful, practical findings (Power et al., 2022; Pearce et al., 2023; Allen-Zhu and Li, 2024; Wang et al., 2024a; Mondorf et al., 2025; Bertolazzi et al., 2025).